# RTG-SLAM: Real-time 3D Reconstruction at Scale Using Gaussian Splatting

### v2

@@ 3.1. Compact Gaussian Representation
The core of optimizing Gaussians lies in rendering color and depth maps through differentiable splatting, calculating errors with input RGBD images, and updating the Gaussian parameters. Now we introduce the rendering process in detail. Given a camera pose $\mathbf{T}_{g}$ and camera intrinsic matrix $\mathbf{K}$, the ray through the center of each pixel $u$ in the image is defined as:
$\mathbf{r}(\mathbf{u})=(\mathbf{R}_{g}\mathbf{K}^{-1}\mathbf{u})\theta+\mathbf{t}_{g},\;\text{where}\,\mathbf{T}_{g}=\left[\begin{array}[]{cc}\mathbf{R}_{g}&\mathbf{t}_{g}\\ 0&1\end{array}\right]\in\mathbb{SE}(3).$
Here $\theta$ is the length parameter along the ray direction. Then the color image $\hat{\mathbf{C}}$ can be rendered by alpha-blending proposed in (Kerbl et al., 2023) :
@@
(2) $f(\mathbf{u})=\alpha_{i}\exp(-\frac{1}{2}(\mathbf{u}-\bm{\mu})^{T}\bm{\Sigma}_{2D,i}^{-1}(\mathbf{u}-\bm{\mu})).$
@@
(4) $\mathbf{p}_{G_{j}^{\mathbf{r}},\mathbf{r}}=(\mathbf{R}_{g}\mathbf{K}^{-1}\mathbf{u})\theta_{\mathbf{u}}+\mathbf{t}_{g},\;\text{where}\,\theta_{\mathbf{u}}=\frac{(\mathbf{p}_{j}^{\mathbf{r}}-\mathbf{t}_{g})\cdot\mathbf{n}_{j}^{\mathbf{r}}}{(\mathbf{R}_{g}\mathbf{K}^{-1}\mathbf{u})\cdot\mathbf{n}_{j}^{\mathbf{r}}}.$
@@ 3.2. Online Reconstruction Process
$\mathbf{G}_{o}=(1-w_{curr})\mathbf{G}_{o-1}+w_{curr}\mathbf{G}^{\prime}_{o},$
@@
We utilize the frame-to-model ICP as the front-end odometry for camera tracking. Specifically, we use the optimized Gaussians in the previous frame to render the depth map $\hat{\mathbf{D}}_{k-1}^{*}$ and normal map $\hat{\mathbf{N}}_{k-1}^{*}$, and convert $\hat{\mathbf{D}}_{k-1}^{*}$ to the global space $\hat{\mathbf{V}}_{k-1}^{g*}$. Then given the current frame $\mathbf{V}_{k}^{l}$, we aim to find the camera pose $\xi$ that minimizes the point-to-plane error between 3D back-projected vertices:
(10) $E(\xi)=\sum\left\|\big{(}\mathbf{T}_{g,k}\mathbf{V}_{k}^{l}(\mathbf{u})-\hat{\mathbf{V}}_{k-1}^{g*}(\hat{u})\big{)}\cdot\hat{\mathbf{N}}_{k-1}^{*}(\hat{u})\right\|.$
$\mathbf{T}_{g,k}=exp(\xi)$ is the matrix exponential that maps a member of the Lie algebra. We run a multi-level ICP to solve the objective function as (Newcombe et al., 2011). Meanwhile, in order to reduce the drift during the scanning of large scenes, we also run a back-end optimization thread similar to ORB-SLAM2 (2017). While the pose estimation is finished, a set of 3D landmarks are also maintained. These landmarks are used for graph optimization in the back-end, enabling more accurate camera tracking.

### v3

@@ 3.1. Compact Gaussian Representation
The core of optimizing Gaussians lies in rendering color and depth maps through differentiable splatting, calculating errors with input RGBD images, and updating the Gaussian parameters. Now we introduce the rendering process in detail. Given a camera pose $\mathbf{T}_{g}$ and camera intrinsic matrix $\mathbf{K}$, the ray through the center of each pixel $\mathbf{u}$ in the image is defined as:
$\mathbf{r}(\mathbf{u})=(\mathbf{R}_{g}\mathbf{K}^{-1}\dot{\mathbf{u}})\theta+\mathbf{t}_{g},\;\text{where}\,\mathbf{T}_{g}=\left[\begin{array}[]{cc}\mathbf{R}_{g}&\mathbf{t}_{g}\\ 0&1\end{array}\right]\in\mathbb{SE}(3).$
Here $\theta$ is the length parameter along the ray direction and $\dot{\mathbf{u}}$ is the homogeneous vector $\dot{\mathbf{u}}:=(\mathbf{u}^{\top}|1)^{\top}$. Then the color image $\hat{\mathbf{C}}$ can be rendered by alpha-blending proposed in (Kerbl et al., 2023) :
@@
(2) $f(\mathbf{u})=\alpha_{i}\exp(-\frac{1}{2}(\mathbf{u}-\bm{\mu})^{\top}\bm{\Sigma}_{2D,i}^{-1}(\mathbf{u}-\bm{\mu})).$
@@
(4) $\mathbf{p}_{G_{j}^{\mathbf{r}},\mathbf{r}}=(\mathbf{R}_{g}\mathbf{K}^{-1}\dot{\mathbf{u}})\theta_{\mathbf{u}}+\mathbf{t}_{g},\;\text{where}\,\theta_{\mathbf{u}}=\frac{(\mathbf{p}_{j}^{\mathbf{r}}-\mathbf{t}_{g})\cdot\mathbf{n}_{j}^{\mathbf{r}}}{(\mathbf{R}_{g}\mathbf{K}^{-1}\dot{\mathbf{u}})\cdot\mathbf{n}_{j}^{\mathbf{r}}}.$
@@ 3.2. Online Reconstruction Process
$G_{o}=(1-w_{curr})G_{o-1}+w_{curr}G^{\prime}_{o},$
@@
We utilize the frame-to-model ICP as the front-end odometry for camera tracking. Specifically, we use the optimized Gaussians in the previous frame to render the depth map $\hat{\mathbf{D}}_{k-1}^{*}$ and normal map $\hat{\mathbf{N}}_{k-1}^{*}$, and convert $\hat{\mathbf{D}}_{k-1}^{*}$ to the global space $\hat{\mathbf{V}}_{k-1}^{g*}$. Then given the current frame $\mathbf{V}_{k}^{l}$, we aim to find the camera pose that minimizes the point-to-plane error between 3D back-projected vertices:
(10) $E(\bm{\xi})=\sum\left\|\big{(}\mathbf{T}_{g,k}\mathbf{V}_{k}^{l}(\mathbf{u})-\hat{\mathbf{V}}_{k-1}^{g*}(\hat{\mathbf{u}})\big{)}\cdot\hat{\mathbf{N}}_{k-1}^{*}(\hat{\mathbf{u}})\right\|.$
Here $\bm{\xi}$ is the Lie algebra representation of $\mathbf{T}_{g,k}$. We run a multi-level ICP to solve the objective function as (Newcombe et al., 2011). Meanwhile, in order to reduce the drift during the scanning of large scenes, we also run a back-end optimization thread similar to ORB-SLAM2 (2017). While the pose estimation is finished, a set of 3D landmarks are also maintained. These landmarks are used for graph optimization in the back-end, enabling more accurate camera tracking.